3,4, and 5 please!
15 points Answer the following questions. Please show your work to help justify your Show critical thinking skills Narrative. J.R is a 27-year-old male. He is 5 feet, 9 inches tall. He weights 187 pounds. He runs three days a week, a moderate pace, for 20 minutes. On the weekends, he enjoys doing yardwork, walking his dog, and hanging out with his friends. J.R. would like to improve his overall diet, and he has the following questions answer 1. [3 points] How many calories per day should J.R. consume? a. Convert weight in pounds to weight in kilograms 187 pounds x 0.454 kilograms / 1 pound 84.898 kg b. Convert height to meters 69 inches x 0.0254 meters 1 inch 1.753 m Use EER equation and PA factor from chapter 9, slides 18-19 to calculate calories per day EER 662-(9.53 x 27)+1.25 [(15.91 x 84.898)+(539.6 x 1.753)] 2769.8 cal/day c. 13 points] How many grams of protein per day should J.R consume? a Convert weight in pounds to weight in kilograms. b Identify how many grams of protein per kilogram of body weight is c. Caiculate total grams of protein per day based on a & b 187 pounds x 0.454 kilograms / 1 pound 84.898 kg appropriate 0.8 g/kg of body weight 0.8 g x 84.898 kg 67.9 3. [3 points] How many grams of carbohydrate per day should J.R. consume? Calculate 45-65% of J Rs calories. This is the range you will need to caiculate grams of carbohydrate per day. Note: use the total calories calculated in question 1 a. b. How many calories per gram are offered from carbohydrates? c. Calculate total grams of carbohydrate per day based on a & b 4. [3 points] How many grams of fat per day should J.R. consume? a. Calculate 20-35% of JRs calories. This is the range you will need to calculate grams of fat per day. Note use the total calories calculated in 898 question 1 b. How many calories per gram are offered from fat? c. Calculate total grams of fat per day based on a & t 5. [3 points] Estimate fluid needs based on a. Energy needs (1 mL per 1 calorie) b. Body weight (25 mL per 1 kg body weight) c. Age, gender (see DRI for fluid/water
Answer #1

3.a. Total calorie intake = 2769.8 cal/day

45-65% carbohydrates=1246- 1800 calories / day.(2769.8*45/100 to 2769.8*65/100)

b.Carbohydrates provide 4 calories per gram.

c.To calculate carbohydrates 45-65% of the calories should be divided by 4.

45-65% = 311 - 450 grams /day carbohydrates should be taken by Mr.J.R

4.a . 20-35%fats = 553- 969 calories /day. (2769.8*20/100 to 2769.8*35/100)

b.Fats provide 9 calories per gram.

c.To calculate fats 20-35% of the calories should be divided by 9.

20-35%=61- 107 grams/ day fat should be consumed by Mr.J.R

5.a.Energy needs 1ml per calorie = 2769.8 ml water is needed.=2.7 litres of fluid is needed to be consumed.

b.Body weight = 25ml /kg =2769.8/25= 3.4 ml of fluid Mr.J.R needs for every calorie Mr.J.R ingests.

c.Age / gender daily fluid need for Mr. J.R is about 2.8 litres and in days he do exercises extra of 0.7 L should be needed.
Answer #2
6,7,8,9,10,11,12,13,14,and15 please!and show work